# Ensembling Portfolio Strategies for Long-Term Investments: A Distribution-Free Preference Framework for Decision-Making and Algorithms

### v1

@@ 2 Model settings and formalizations
Following the portfolios selected at any time $n$, denoted as $b_{n}\in\mathcal{B}^{m}$, let us denote a strategy associated with them by the infinite sequence $\left(b_{n}\right)\coloneqq\left\{b_{n}\right\}_{n=1}^{\infty}\in\mathcal{B}^{m\times\infty}$; accordingly, a constant strategy of a fixed portfolio $b$ is denoted as solely $\left(b\right)$. Given the initial capital $S_{0}=1$, assuming there are no commission fees, the cumulative wealth and its corresponding exponential average growth rate after $n$ periods of investment by a generic strategy $\left(b_{n}\right)$ are defined respectively as:
@@
Figure 1 illustrates an example for the evolutions of some strategies taken from the experiment section. The left graphic shows two strategies oftenly exceeding each other while the right graphic shows a strategy almost always exceeding the remainder, except for a few observations around time $n=3400$, where they are temporarily equivalent.
Refer to captionRefer to caption
@@ 2.1 Preference relation and combination criterion
Given a set of observable strategies, since the infinite sequence $x_{1}^{\infty}$ is unknown, their evolutions of cumulative wealths are not definitively determined at any time. Thus, any competitive strategy can only guarantee to be at least as good as the best strategy at some future particular observation periods. In order to create such a competitive strategy conveniently and comprehensively, it is necessary to define a single baseline strategy against which it will compete. This baseline strategy, representing the best possible cumulative wealth among the strategies observed, is formulated based on the realizations of assets’ returns during the respective time periods.
@@
(Baseline strategy). Given an infinite sequence of assets’ returns $x_{1}^{\infty}$ and a set of strategies $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\coloneqq\big{\{}1,\ldots,k\big{\}}\big{\}}$, the baseline strategy, denoted as $\big{(}\ddot{b}_{n}\big{)}$ hereafter, is the best merged strategy of all the component ones as $S_{n}\big{(}\ddot{b}_{n}\big{)}\coloneqq\max_{\alpha\in[k]}S_{n}\big{(}b_{n}^{\alpha}\big{)}$ for all $n$.
By definition, the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$ is the envelope of all cumulative wealth of observed strategies over time, so $\big{(}\ddot{b}_{n}\big{)}\succsim\big{(}b_{n}^{\alpha}\big{)}$ for all $\alpha\in\left[k\right]$. Since the baseline strategy’s construction is done using future data of assets’ returns at each time period, it is basically inaccessible for the investor; however, the investor could create a multi-strategy (or an ensemble) instead, denoted as $\big{(}b_{n}\big{)}$ and termed the combinatorial strategy henceforth, to compete with the baseline strategy. Despite the inaccessibility of the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$, it facilitates us to evaluate the preference between the combinatorial strategy and a single consistent strategy in the infinite-dimensional space $\mathcal{B}^{m\times\infty}$ so that the investor’s objective could be guaranteed. To clarify this point, it is simple but worth reminding that if $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$, then $\big{(}b_{n}\big{)}\succ\big{(}b_{n}^{\alpha}\big{)}$ for all $\alpha\in\left[k\right]$ due to the transitivity and asymmetry of preference, but the reverse is not necessarily true. Since the strict preference is favored than the weak one, let us establish a specific criterion for combining the strategies.
@@
Consider two generic strategies $\big{(}\hat{b}_{n}\big{)}$ and $\big{(}\bar{b}_{n}\big{)}$ and the infinite-dimensional space $\mathbb{R}_{+}^{m\times\infty}$ of all instances of the sequence $x_{1}^{\infty}$. By the completeness and asymmetry of preference, if $\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}$ everywhere in $\mathbb{R}_{+}^{m\times\infty}$ while $\big{(}\hat{b}_{n}\big{)}\succsim\big{(}\bar{b}_{n}\big{)}$ only in a non-empty set $D\subset\mathbb{R}_{+}^{m\times\infty}$, then $\big{(}\bar{b}_{n}\big{)}\succ\big{(}\hat{b}_{n}\big{)}$ in $\mathbb{R}_{+}^{m\times\infty}\backslash D$. In this context, an economically rational investor in the market universally prefers the strategy $\big{(}\bar{b}_{n}\big{)}$ over the strategy $\big{(}\hat{b}_{n}\big{)}$, or the preference $\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}$ is said to be more universal than the preference $\big{(}\hat{b}_{n}\big{)}\succsim\big{(}\bar{b}_{n}\big{)}$. This is reasonable by the fact that the possibility of the strategy $\big{(}\hat{b}_{n}\big{)}$ being strictly preferred does not exist. In the case $\big{(}\bar{b}_{n}\big{)}\sim\big{(}\hat{b}_{n}\big{)}$ everywhere in $\mathbb{R}_{+}^{m\times\infty}$, they are said to be universally indifferent.
@@ 3.1 The benchmark strategy
In this section, a benchmark strategy, denoted as $\big{(}b_{n}^{*}\big{)}$, is formed using the concept in Definition 4 such that the preference $\big{(}b_{n}^{*}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ is more universal than the reverse one $\big{(}\ddot{b}_{n}\big{)}\succsim\big{(}b_{n}^{*}\big{)}$ in the sense of Claim 1. Although the strategy $\big{(}b_{n}^{*}\big{)}$ is unattainable as it relies on future data like the baseline $\big{(}\ddot{b}_{n}\big{)}$, we could create a combinatorial strategy $\big{(}b_{n}\big{)}$ using only past data such that the preference $\big{(}b_{n}\big{)}\sim\big{(}b_{n}^{*}\big{)}$ holds everywhere in $\mathbb{R}_{+}^{m\times\infty}$, making $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ more universal than the reverse preference by the transitivity of the preference relation.
@@
Definition 5 immediately results in $W_{n}\big{(}b_{n}^{*}\big{)}\geq W_{n}\big{(}\ddot{b}_{n}\big{)}$, so the preference $\big{(}b_{n}^{*}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ holds for any sequence $x_{1}^{\infty}$. If $\big{(}b_{n}^{*}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ for some sequences in $\mathbb{R}_{+}^{m\times\infty}$, any combinatorial strategy satisfying $\big{(}b_{n}\big{)}\sim\big{(}b_{n}^{*}\big{)}$ implies $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ due to the transitivity and asymmetry of the preference relation, thus meeting the Combination criterion. Therefore, it is sufficient to secure the preference $\big{(}b_{n}^{*}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ for a sequence $x_{1}^{\infty}$ by verifying if $\big{(}W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}\ddot{b}_{n}\big{)}\big{)}\nrightarrow 0$ as $n\to\infty$ due to $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}\land\big{(}b_{n}\big{)}\nsim\big{(}\ddot{b}_{n}\big{)}\Leftrightarrow\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$.
@@ 3.2 Small scale of component strategies
In this section, we consider the case of a small number $k\geq 2$ for the set $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$ of component strategies. After introducing a specific benchmark strategy $\big{(}b_{n}^{*}\big{)}$ in the previous section, we propose now a combinatorial strategy $\big{(}b_{n}\big{)}$, which does not depend on future data, such that $\big{(}b_{n}\big{)}\sim\big{(}b_{n}^{*}\big{)}$ for all instances of sequences in $\mathbb{R}_{+}^{m\times\infty}$. This goal can be accomplished by demonstrating that the distance $\left|\log S_{n}\big{(}b_{n}^{*}\big{)}-\log S_{n}\big{(}b_{n}\big{)}\right|$ in the worst case of sequence $x_{1}^{n}$ is sublinear in the time variable $n$. To this end, we propose a combinatorial strategy as follows:
@@
In this case, the combinatorial strategy is equivalent to the exponentially weighted average forecaster with the learning rate parameter of one and the logarithmic loss function, as established in Cesa-Bianchi et al. (1997), Cesa-Bianchi and Lugosi (1999, 2006). Proposition 4, given below, establishes an upper bound on the positive difference $\log S_{n}\big{(}b_{n}^{*}\big{)}-\log S_{n}\big{(}b_{n}\big{)}$ in the worst case of the sequence $x_{1}^{n}$. It follows from the theorem for the minimax problem shown in Cover (1991), Cover and Ordentlich (1996), Cover and Thomas (2006) and Cesa-Bianchi and Lugosi (2006) [^2] Within the framework of Prediction with expert advice, as discussed in Cesa-Bianchi et al. (1997), Cesa-Bianchi and Lugosi (1999, 2006), the Universal portfolio is demonstrated to be equivalent to the Laplace mixture forecaster, studied in Davisson (1973), Rissanen (1986), for the case of uniform distribution, and the Krichevsky–Trofimov mixture forecaster, proposed by Krichevsky and Trofimov (1981), for the Dirichlet distribution in the Kelly betting market. It has also been investigated within another umbrella of Online portfolio selection, as in Ordentlich and Cover (1996), Helmbold et al. (1998), Blum and Kalai (1999), Fostera and Vohra (1999), Vovk and Watkins (1998), van Erven et al. (2020)..
@@ 3.3 Arbitrarily large scale of component strategies
This section addresses the guarantee of the proposed combinatorial strategy in the previous section when the scale of the set $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$ may be exceedingly large, even infinite. In general, the limit ${\displaystyle\left({\textstyle W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}b_{n}\big{)}}\right)}\to 0$ is not guaranteed because the upper bound established in Proposition 4 is linear in $k$, meaning that $\left(k-1\right)\log\left(n+1\right)\to\infty$ as $k\to\infty$. Furthermore, when $k\to\infty$, a convex combination of all the component portfolios results in atomic weights, rendering the combination meaningless. As a result, the combinatorial strategy becomes nothing more than an exponentially weighted average of all component strategies, as formed by the formula (4.1) in the experiment section, failing to meet the Combination criterion as $\big{(}b_{n}^{*}\big{)}\sim\big{(}\ddot{b}_{n}\big{)}$ for any sequence $x_{1}^{\infty}$. To address this issue, we propose a modified combinatorial strategy $\big{(}b_{n}\big{)}$ below.
@@
$\hat{b}_{1}^{i}\coloneqq{\displaystyle\sum_{\mathcal{A}^{i}}{\displaystyle b_{1}^{\alpha}}}\mu^{i}\big{(}\alpha\big{)},\,\forall i\in\left[N\right]\text{ and }\hat{b}_{n}^{i}\coloneqq\dfrac{{\displaystyle\sum_{\mathcal{A}^{i}}{\displaystyle b_{n}^{\alpha}}}S_{n-1}\left(b_{n-1}^{\alpha}\right)\mu^{i}\left(\alpha\right)}{{\displaystyle\sum_{\mathcal{A}^{i}}}{\displaystyle S_{n-1}\left(b_{n-1}^{\alpha}\right)\mu^{i}\left(\alpha\right)}},\forall n\geq 2,i\in\left[N\right],$ (3.4)
where each $\mu^{i}\left(\alpha\right)$ is the probability mass with respect to the variable $\alpha\in\mathcal{A}^{i}$ and $i\in\left[N\right]$.
The proposed combinatorial strategy utilizes the partition to reduce dimensionality, where each base set represents a population of component strategies corresponding to a distribution. Particularly, the representative strategies $\big{(}\hat{b}_{n}^{i}\big{)}$, as per (3.4), represent aggregate behaviors of all the component strategies in each corresponding base set with respect to the mass $\mu^{i}\left(\alpha\right)$; then the combinatorial strategy $\big{(}b_{n}\big{)}$ is constructed based on these representative strategies according to (3.3), which is similar to the formula (3.1). In addition, concerning the allocation of wealth to the component strategies, let $P_{n}\big{(}\alpha\big{)}$ denote the time-variant distribution of wealth on the component strategies, and consider the collection of these distributions as a multi-strategy. Then, for any $n$, the capital allocated to each component portfolio $b_{n}^{\alpha}$ is as:
$P_{n}\big{(}\alpha\big{)}={\displaystyle\dfrac{{\displaystyle\int_{\mathcal{B}^{k}}{\displaystyle\lambda_{i}}}S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}S_{n-1}\big{(}b_{n-1}^{\alpha}\big{)}\mu^{i}\big{(}\alpha\big{)}\mu\left(\lambda\right)\mathop{}\!\mathrm{d}\lambda}{{\displaystyle\int_{\mathcal{B}^{k}}\sum_{\mathcal{A}^{i}}S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}{\displaystyle S_{n-1}\big{(}b_{n-1}^{\alpha}\big{)}\mu^{i}\left(\alpha\right)\mu\left(\lambda\right)}}\mathop{}\!\mathrm{d}\lambda}},\,\forall\alpha\in\mathcal{A}^{i},i\in\left[N\right],$
so the return at time $n$ of the combinatorial portfolio $\big{(}b_{n}\big{)}$ is ${\displaystyle\sum}_{\mathcal{A}}\left\langle b_{n}^{\alpha},x_{n}\right\rangle P_{n}\big{(}\alpha\big{)}$.
@@
If the minimal mass values $\epsilon_{n}\gg 0$ for any $n\geq 2$, then by Proposition 5, we have $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ for any sequence $x_{1}^{\infty}$ as the upper bound converges to zero. The representative strategies with uniform probability mass $\mu^{i}\left(\alpha\right)$ are instances that ensure $\epsilon_{n}\gg 0$ over the countable base sets. As we noted above, the infimum limit is not bounded by zero, so there exist sequences $x_{1}^{\infty}$ where the combinatorial strategy $\big{(}b_{n}\big{)}$ is not always exceeded by the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$. The remark below specifies conditions to guarantee the Combination criterion. In addition, the time required for the combinatorial strategy to exceed all component strategies increases significantly with larger numbers of base sets and their respective scales. Additionally, the partitioning into base sets affects the condition $\big{(}W_{n}\big{(}b_{n}\big{)}-W_{n}\big{(}\dot{b}_{t}\big{)}\big{)}\nrightarrow 0$ and the evolutions of representative strategies’ cumulative wealths. Furthermore, the combinatorial strategy reduces to the one proposed for a small number $k$ of component strategies by equating $N=k$.
@@
Let $\big{(}\dot{b}_{n}\big{)}$ denote the best merged strategy of all the representative strategies, as similarly defined as the baseline strategy in Definition 3. The benchmark strategy $\big{(}b_{n}^{*}\big{)}$ is defined as the constant combination of the representative strategies, not the component strategies. If $n^{-1}\log\epsilon_{n}\to 0$, then $\big{(}\ddot{b}_{n}\big{)}\sim\big{(}\dot{b}_{n}\big{)}$ for any sequence $x_{1}^{\infty}$; moreover, for some sequence such that $\big{(}W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}\dot{b}_{n}\big{)}\big{)}\nrightarrow 0$, then $\big{(}b_{n}\big{)}\succ\big{(}\dot{b}_{n}\big{)}$ so $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ which satisfies the Combination criterion, due to the transitivity and asymmetry of the preference relation.

### v2

@@ 2 Model settings and formalizations
At each $n$, let $b_{n}:\mathbb{R}_{++}^{m\times(n-1)}\to\mathcal{B}^{m}$ denote the choice of portfolio based on $x_{1}^{n-1}$, and let the strategy associated with these portfolios be denoted by the infinite sequence $\big{(}b_{n}\big{)}\coloneqq\left\{b_{n}\right\}_{n=1}^{\infty}\in\mathcal{B}^{m\times\infty}$; accordingly, a constant strategy of a fixed portfolio $b$ is denoted as solely $\big{(}b\big{)}$, without a timing index. Given the initial capital $S_{0}\big{(}b_{0}\big{)}\eqqcolon S_{0}=1$, assuming there are no commission fees, the cumulative wealth and its corresponding exponential average growth rate after $n$ periods of investment using a generic strategy $\left(b_{n}\right)$ are defined respectively as:
@@
Figure 1 illustrates an example of the evolutions of some strategies taken from the experiment section. The left graphic shows two strategies often exceeding each other, while the right graphic shows a strategy almost always exceeding the remainder, except for a few observations around time $n=3400$, where they are temporarily equivalent.
@@ 2.1 Preference relation and combination criterion
Given a set of strategies, since the infinite sequence $x_{1}^{\infty}$ is unknown, their evolutions of cumulative wealths are not definitively determined at any time. Thus, any competitive strategy can only guarantee to be at least as good as the best strategy at some future particular observation periods. In order to create such a competitive strategy conveniently and comprehensively, it is necessary to define a single baseline strategy against which it will compete. This baseline strategy, representing the best possible cumulative wealth among the observed strategies, is formulated based on the realizations of assets’ returns during the respective time periods.
@@
(Baseline strategy). For an infinite sequence of assets’ returns $x_{1}^{\infty}$ and a set of strategies $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\coloneqq\big{\{}1,\ldots,k\big{\}}\big{\}}$, the baseline strategy, denoted as $\big{(}\ddot{b}_{n}\big{)}$ hereafter, is defined as the best merged strategy of all the component ones, where $S_{n}\big{(}\ddot{b}_{n}\big{)}\coloneqq\max_{\alpha\in[k]}S_{n}\big{(}b_{n}^{\alpha}\big{)}$ for all $n$.
By definition, the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$ is the envelope of all cumulative wealth of observed strategies over time, so $\big{(}\ddot{b}_{n}\big{)}\succsim\big{(}b_{n}^{\alpha}\big{)}$ for all $\alpha\in\left[k\right]$. Since the construction of the baseline strategy relies on future data of assets’ returns at every time period, it is inherently inaccessible to the investor. However, the investor could instead create a multi-strategy (or an ensemble), denoted as $\big{(}b_{n}\big{)}$ and termed the combinatorial strategy henceforth, to compete with the baseline strategy. Despite the inaccessibility of the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$, it facilitates us to evaluate the preference between the combinatorial strategy and a single consistent strategy in the infinite-dimensional space $\mathcal{B}^{m\times\infty}$, thereby ensuring that the investor’s objective is met. To clarify this point, it is simple but worth noting that if $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$, then $\big{(}b_{n}\big{)}\succ\big{(}b_{n}^{\alpha}\big{)}$ for all $\alpha\in\left[k\right]$, due to the transitivity and asymmetry of preference, but the reverse is not necessarily true. Since strict preference is favored over a weak one, let us establish a specific criterion for combining the strategies.
@@
Consider two generic strategies $\big{(}\hat{b}_{n}\big{)}$ and $\big{(}\bar{b}_{n}\big{)}$ and the infinite-dimensional distribution space $\mathbb{R}_{++}^{m\times\infty}$ of all instances of the sequence $x_{1}^{\infty}$. By the completeness and asymmetry of preference, if $\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}$ everywhere in $\mathbb{R}_{++}^{m\times\infty}$ while $\big{(}\hat{b}_{n}\big{)}\succsim\big{(}\bar{b}_{n}\big{)}$ only in a non-empty set $D\subset\mathbb{R}_{++}^{m\times\infty}$, then $\big{(}\bar{b}_{n}\big{)}\succ\big{(}\hat{b}_{n}\big{)}$ in $\mathbb{R}_{++}^{m\times\infty}\backslash D$. In this context, an economically rational investor in the market universally prefers the strategy $\big{(}\bar{b}_{n}\big{)}$ over the strategy $\big{(}\hat{b}_{n}\big{)}$, or the preference $\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}$ is said to be more universal than the preference $\big{(}\hat{b}_{n}\big{)}\succsim\big{(}\bar{b}_{n}\big{)}$. This is reasonable by the fact that the possibility of the strategy $\big{(}\hat{b}_{n}\big{)}$ being strictly preferred does not exist. In the case $\big{(}\bar{b}_{n}\big{)}\sim\big{(}\hat{b}_{n}\big{)}$ everywhere in $\mathbb{R}_{++}^{m\times\infty}$, they are said to be universally indifferent.
@@ 3.1 The benchmark strategy
In this section, a benchmark strategy, denoted as $\big{(}b_{n}^{*}\big{)}$, is formed using the concept in Definition 4 such that the preference $\big{(}b_{n}^{*}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ is more universal than the reverse one $\big{(}\ddot{b}_{n}\big{)}\succsim\big{(}b_{n}^{*}\big{)}$ in the sense of Claim 1. Although the strategy $\big{(}b_{n}^{*}\big{)}$ is unattainable as it relies on future data like the baseline $\big{(}\ddot{b}_{n}\big{)}$, we could create a combinatorial strategy $\big{(}b_{n}\big{)}$ using only past data such that the preference $\big{(}b_{n}\big{)}\sim\big{(}b_{n}^{*}\big{)}$ holds everywhere in $\mathbb{R}_{++}^{m\times\infty}$, making $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ more universal than the reverse preference by the transitivity of the preference relation.
@@
Definition 5 immediately results in $W_{n}\big{(}b_{n}^{*}\big{)}\geq W_{n}\big{(}\ddot{b}_{n}\big{)}$, so the preference $\big{(}b_{n}^{*}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ holds for any sequence $x_{1}^{\infty}$. If $\big{(}b_{n}^{*}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ for some sequences in $\mathbb{R}_{++}^{m\times\infty}$, any combinatorial strategy satisfying $\big{(}b_{n}\big{)}\sim\big{(}b_{n}^{*}\big{)}$ implies $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ due to the transitivity and asymmetry of the preference relation, thus meeting the Combination criterion. Therefore, it is sufficient to secure the preference $\big{(}b_{n}^{*}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ for a sequence $x_{1}^{\infty}$ by verifying if $\big{(}W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}\ddot{b}_{n}\big{)}\big{)}\nrightarrow 0$ as $n\to\infty$ due to $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}\land\big{(}b_{n}\big{)}\nsim\big{(}\ddot{b}_{n}\big{)}\Leftrightarrow\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$.
@@ 3.2 Small scale of component strategies
In this section, we consider the case of a small number $k\geq 2$ for the set $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$ of component strategies. After introducing a specific benchmark strategy $\big{(}b_{n}^{*}\big{)}$ in the previous section, we propose now a combinatorial strategy $\big{(}b_{n}\big{)}$, which does not depend on future data, such that $\big{(}b_{n}\big{)}\sim\big{(}b_{n}^{*}\big{)}$ for all instances of sequences in $\mathbb{R}_{++}^{m\times\infty}$. This goal can be accomplished by demonstrating that the distance $\left|\log S_{n}\big{(}b_{n}^{*}\big{)}-\log S_{n}\big{(}b_{n}\big{)}\right|$ in the worst case of sequence $x_{1}^{n}$ is sublinear in the time variable $n$. To this end, we propose a combinatorial strategy as follows:
@@
In this case, the combinatorial strategy is equivalent to the exponentially weighted average forecaster with the learning rate parameter of one and the logarithmic loss function, as established in Cesa-Bianchi et al. (1997), Cesa-Bianchi and Lugosi (1999, 2006). Proposition 4, given below, establishes an upper bound on the positive difference $\log S_{n}\big{(}b_{n}^{*}\big{)}-\log S_{n}\big{(}b_{n}\big{)}$ in the worst case of the sequence $x_{1}^{n}$. It follows from the theorem for the minimax problem shown in Cover (1991), Cover and Ordentlich (1996), Cover and Thomas (2006) and Cesa-Bianchi and Lugosi (2006) [^2] Within the framework of Prediction with expert advice, as discussed in Cesa-Bianchi et al. (1997), Cesa-Bianchi and Lugosi (1999, 2006), the Universal portfolio is equivalent to the Laplace mixture forecaster, studied in Davisson (1973), Rissanen (1986), for the case of uniform distribution, and to the Krichevsky–Trofimov mixture forecaster, proposed by Krichevsky and Trofimov (1981), for the Dirichlet distribution in the Kelly betting market. It has also been investigated within another umbrella of Online portfolio selection, as in Ordentlich and Cover (1996), Helmbold et al. (1998), Blum and Kalai (1999), Fostera and Vohra (1999), Vovk and Watkins (1998), van Erven et al. (2020)..
@@ 3.3 Arbitrarily large scale of component strategies
When the scale of the set $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$ becomes larger, the computation of the proposed combinatorial strategy becomes increasingly difficult, especially at relatively large scales. Furthermore, the limit ${\displaystyle\big{(}{\textstyle W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}b_{n}\big{)}}\big{)}}\to 0$ is generally not guaranteed as $k\to\infty$, since the upper bound established in Proposition 4 is linear in $k$, meaning that $\left(k-1\right)\log\left(n+1\right)\to\infty$. Additionally, when $k\to\infty$, a convex combination of all the component portfolios results in atomic weights, rendering the combination meaningless. As a result, the combinatorial strategy reduces to nothing more than an exponentially weighted average of all component strategies, as formed by the formula (4.1) in the experiment section, thereby failing to meet the Combination criterion, as $\big{(}b_{n}^{*}\big{)}\sim\big{(}\ddot{b}_{n}\big{)}$ for any sequence $x_{1}^{\infty}$. To address this issue, we propose a modified combinatorial strategy below.
@@
$\hat{b}_{1}^{i}\coloneqq{\displaystyle\sum_{\alpha\in\mathcal{A}^{i}}{\displaystyle b_{1}^{\alpha}}}\mu^{i}\big{(}\alpha\big{)},\,\forall i\in\left[N\right]\text{ and }\hat{b}_{n}^{i}\coloneqq\dfrac{{\displaystyle{\displaystyle\sum_{\alpha\in\mathcal{A}^{i}}}{\displaystyle b_{n}^{\alpha}}}S_{n-1}\left(b_{n-1}^{\alpha}\right)\mu^{i}\left(\alpha\right)}{{\displaystyle\sum_{\alpha\in\mathcal{A}^{i}}}{\displaystyle S_{n-1}\left(b_{n-1}^{\alpha}\right)\mu^{i}\left(\alpha\right)}},\forall n\geq 2,i\in\left[N\right],$ (3.4)
where each $\mu^{i}\left(\alpha\right)$ is the probability mass with respect to $\alpha\in\mathcal{A}^{i}$ and $i\in\left[N\right]$.
The proposed combinatorial strategy utilizes the partition to reduce dimensionality, where each base set represents a population of component strategies corresponding to a distribution. Particularly, the representative strategies $\big{(}\hat{b}_{n}^{i}\big{)}$, as per (3.4), represent aggregate behaviors of all the component strategies in each corresponding base set with respect to the mass $\mu^{i}\left(\alpha\right)$; then the combinatorial strategy $\big{(}b_{n}\big{)}$ is constructed based on these representative strategies according to (3.3), which is similar to the formula (3.1). In addition, concerning the allocation of wealth among the component strategies, let $P_{n}\big{(}b_{n}^{\alpha}\big{)}$ denote the time-dependent distribution of wealth over the corresponding component strategies $\big{(}b_{n}^{\alpha}\big{)}$; then, the capital allocated to each component portfolio $b_{n}^{\alpha}$ at time $n$ is given as follows:
$P_{n}\big{(}b_{n}^{\alpha}\big{)}={\displaystyle\dfrac{{\displaystyle\int_{\mathcal{B}^{k}}{\displaystyle\lambda_{i}}}S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}S_{n-1}\big{(}b_{n-1}^{\alpha}\big{)}\mu^{i}\big{(}\alpha\big{)}\mu\left(\lambda\right)\mathop{}\!\mathrm{d}\lambda}{{\displaystyle\int_{\mathcal{B}^{k}}\sum_{\beta\in\mathcal{A}^{i}}S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}{\displaystyle S_{n-1}\big{(}b_{n-1}^{\beta}\big{)}\mu^{i}\left(\beta\right)\mu\left(\lambda\right)}}\mathop{}\!\mathrm{d}\lambda}},\,\forall\alpha\in\mathcal{A}^{i},i\in\left[N\right].$
Thus, the return at time $n$ of the combinatorial portfolio $\big{(}b_{n}\big{)}$ is ${\displaystyle\sum}_{\alpha\in\left[k\right]}\left\langle b_{n}^{\alpha},x_{n}\right\rangle P_{n}\big{(}\alpha\big{)}$.
@@
If the minimal mass values $\epsilon_{n}$ are chosen such that $n^{-1}\log\epsilon_{n}\to 0$, then by Proposition 5, we have $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ for any sequence $x_{1}^{\infty}$, as the upper bound converges to zero. The case of representative strategies with uniform probability mass $\mu^{i}\left(\alpha\right)$ over each countable base set is an instance ensuring such convergence. As we noted above, the infimum limit is not bounded by zero, so there exist sequences $x_{1}^{\infty}$ where the combinatorial strategy $\big{(}b_{n}\big{)}$ is not always exceeded by the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$. The remark below specifies conditions to guarantee the Combination criterion. In addition, the time required for the combinatorial strategy to exceed all component strategies increases significantly with larger numbers of base sets and their respective scales. Additionally, the partitioning into base sets affects the condition $\big{(}W_{n}\big{(}b_{n}\big{)}-W_{n}\big{(}\dot{b}_{t}\big{)}\big{)}\nrightarrow 0$ and the evolutions of representative strategies’ cumulative wealths. Furthermore, the combinatorial strategy reduces to the one proposed for a small number $k$ of component strategies by equating $N=k$.
@@
Let $\big{(}\dot{b}_{n}\big{)}$ denote the best merged strategy of all the representative strategies, as similarly defined as the baseline strategy in Definition 3. The benchmark strategy $\big{(}b_{n}^{*}\big{)}$ is defined as the constant combination of the representative strategies, not the component strategies. If $n^{-1}\log\epsilon_{n}\to 0$, then $\big{(}\ddot{b}_{n}\big{)}\sim\big{(}\dot{b}_{n}\big{)}$ for any sequence $x_{1}^{\infty}$; moreover, for sequences such that $\big{(}W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}\dot{b}_{n}\big{)}\big{)}\nrightarrow 0$, then $\big{(}b_{n}\big{)}\succ\big{(}\dot{b}_{n}\big{)}$ so $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ which satisfies the Combination criterion, due to the transitivity and asymmetry of the preference relation.